# Generative Anchored Fields: Controlled Data Generation via Emergent Velocity Fields and Transport Algebra

### v1

@@ 3.3 Transport Algebra
Transport Algebra organizes GAF’s endpoint pairs into an algebra of emergent velocity fields, enabling composable operations that yield controlled data generation. To understand this, recall that GAF defines two time-conditioned endpoints $J$ and $K$. The velocity field that arises from their disagreement $v(\mathbf{x}_{t},t)=K-J$ defines a time-conditioned vector field. Because GAF is modular, the shared trunk can host multiple tuplets $\{(J_{n},K_{n})\}_{n=1}^{N}$, yielding $v_{n}=K_{n}-J_{n}$ emergent velocity fields. This creates a natural algebra of transport in which vector field operations on $J_{n},K_{n},v_{n}$ correspond to vector combinations of directions in the field. This ability of GAF to operate algebraically leads to a precise and semantically coherent method of data generation across multiple classes and domains.
@@
GAF’s independent $K$ heads enable compositional transport through linear velocity operations. Let $i,j\in\{1,\dots,N\}$.
@@
    Single-class transport. If the noise endpoint $J$ is shared, then
@@
    Cross-modality transport (via $J$). Move between data endpoints of two (or more) domains. To move from class $i$ to class $j$, we use $J$ as a terminal and $v_{i\to j}=K_{i}-K_{j}$ expands to an encode-decode process:
@@
    This sequential transport process uses $J$ as a universal intermediate representation shared across all classes, enabling transport between any pair of $K$ heads without requiring direct $K_{i}\leftrightarrow K_{j}$.
@@
    Direct $K$-head interpolation. We can interpolate directly in the velocity space via weighted combinations
    $v_{\alpha}(\mathbf{x}_{t},t)=(1-\alpha)(K_{i}-J)+\alpha(K_{j}-J)=(1-\alpha)v_{i}+\alpha v_{j},$ (28)
@@
    A simultaneous weighted combination of multiple $K$ heads enables compositional generation, which is experimented on in Section 5.5.3. These operations demonstrate GAF’s transport algebra where $K$ independent heads act as linear operators that can be composed, interpolated, and combined to traverse the generative space. This property extends naturally to sequential domains. For example, for video generation, we hypothesize that GAF can smoothly transition between temporal states while preserving semantic coherence through velocity interpolation, which we will explore in future work.
@@
    Multi-class composition (barycentric). Combine multiple velocity fields by linear vector operations.
    $v_{\text{blend}}=\sum_{m}w_{m}(K_{m}-J),\text{with}\sum_{m}w_{m}=1$
By using $J_{n},K_{n},v_{n}$ that live in $\mathcal{S}$ and the $\{(J_{n},K_{n})\}_{n=1}^{N}$ pairings, one can traverse modalities, switch modalities mid generation, revert to the previous modality, chain back to the origin, or even use a specific generated modality as a bootstrap for a chained generation. All of this is possible by switching between the GAF’s velocity field $v_{i}$, $i\in\{1,\dots,N\}$, and applying simple vector operations. In all cases, integration follows the same ODE solver over the chosen $v$ while the $\mathcal{S}$ operator reconfigures bridges and directions on the fly.
@@ 4.1 Network Architecture
Figure 5: GAF architecture. The DiT-Trunk block follows the architecture of [williampeebles2023].
@@
As shown in Figure 5, we employ a Diffusion Transformer (DiT) [williampeebles2023] architecture as our feature extractor. The GAF trunk takes in either pixel-space representation (e.g., $32\times 32\times 3$ px for low-resolution images) or a latent-space representation (e.g., $32\times 32\times 4$ for VAE, encoded from images with higher resolutions, such as $256\times 256\times 3$ px).
@@
Patchify is the first layer of the trunk, converting the pixel- or latent-space inputs into a sequence of $T$ tokens. It does this by dividing the input in patches, and linearly embedding each patch into the trunk’s hidden dimension $D$. Standard sinusoidal positional embeddings are then added to the tokens to retain spatial information. As noted in [williampeebles2023], the choice of patch size $p$ presents a crucial tradeoff between computational cost and model granularity, as halving $p$ quadruples the GFLOPS with negligible effect on model parameter count. Therefore, we explore $p$ as $2$ and $4$ in our configurations.
Transformer Block.
@@
The $J$ head is a two-layer $\text{MLP}_{J}$, which maps trunk features to noise boundary residuals. The MLP expands from $D$ to $2D$ hidden units via a linear layer, applies GELU activation, and then projects to $p^{2}\times C$ features, where $C$ is the number of channels ($4$ for latents, 3 for RGB pixels). The final layer of $\text{MLP}_{J}$ is initialized to zero for stability. Finally, the outputs are reshaped (i.e., unpatchified) back to their spatial form ($C\times H\times W$).
@@
For each class $k\in\{0,\dots,N-1\}$, each corresponding $K$ head is an independent two-layer $\text{MLP}_{K}$ with identical architecture to $J$. The $K$ heads share no parameters across classes, enabling independent transport operators for each class. During the forward pass, trunk features are computed once for the entire batch, then routed to the appropriate $K$ heads based on class labels. The routing ensures each $K$ head trains on only its assigned class, while the trunk learns shared representations across all classes. Notably, while we report results using this two-layer configuration, we empirically observed that a single linear projection is sufficient for high-fidelity generation, confirming that the backbone $\Phi$ captures the necessary feature information for the twins.
@@ 4.2 Training
We train GAF using three loss components as defined in (24) with mixed-precision training (bfloat16):
@@
where $\mathcal{L}_{\text{pair}}$ anchors the twins $J$ and $K$ to their endpoints ($\mathbf{z}_{y}$ and $\mathbf{z}_{x}$), $\mathcal{L}_{\text{res}}$ penalizes the residuals near endpoints, and $\mathcal{L}_{\text{swap}}$ enforces time-reversal symmetry at the midpoint for consistency and self-correction. Throughout all experiments, we set $\lambda_{\text{res}}=0.003$ and $\lambda_{\text{swap}}=0.002$ during training.
@@ 4.3 Sampling
At inference, we generate samples by integrating the emergent velocity field v=KJv=K-J from noise ($t=0$) to data ($t=1$) using Euler’s method [10.5555/3600270.3602196]. Unlike standard models, GAF defines this field algebraically, enabling precise compositional generation directly within the ODE solver. As during training, we discretize the time interval $t_{\text{eps}},1-t_{\text{eps}}$] into uniformly spaced steps:
$t_{0}=t_{\text{eps}},t_{1},\dots,1-t_{\text{eps}},with~t_{\text{eps}}=1\times 10^{-3}$
@@
For example, in the experiments in Section 5, we use $N=250$ Euler steps for high-quality evaluation. This translates to $250$ Number of Function Evaluations (NFE), as Eleur requires a single model evaluations per step. Additionally, to demonstrate the fast sampling capability, we also report results at $N=20$ (or NFE=$20$).
@@ 5 Experimental Evaluation
In this section, we evaluate GAF for image generation, as proposed in Section 4. We use $N=20$ Euler steps for all generations, with the exception of section 5.2, where we vary the number of steps.
@@ 5.1.1 Dataset
The proposed GAF model was trained on three datasets spanning different visual domains and classes.
CelebA-HQ [liu2015faceattributes] consists of $202,599$ celebrity face images resized to $64\times 64$ px, $256\times 256$ px, and $512\times 512$ px resolutions. We use this dataset for a single-class (i.e., one $J$ and one $K$) generation task.
AFHQ [choi2020starganv2] contains $15,000$ high-quality images of cats, dogs, and wildlife at $512\times 512$ px resolution. We use this dataset for a $3$-class (i.e., one $J$ and three $K$) generation task.
ImageNet-1k [deng2009imagenet, ILSVRC15] is the full ImageNet classification dataset with $1.28$ million training images across $1,000$ object classes, with $1,300$ images per classes. For demonstration purposes, we use two subsets: a single-class dataset (School Bus) and a multi-class subset consisting of $10$ distinct classes. For multi-class generation, we use $10$ $K$ heads, each corresponding to a specific class.
@@ 5.1.2 Training
We train CelebA-64 directly in pixel space, as the images are low-resolution. The other datasets contain higher-resolution images and are therefore first encoded by the SDXL-VAE [podell2024sdxl]. We use the AdamW optimizer with $\beta=(0.9,0.99)$, and learning rate $1\times 10^{-4}$ for all of the datasets, with no weight decay in any of the training runs. The ImageNet-1k bus class is trained for $300,000$ iterations, whereas the other models are trained for $500,000$ iterations.
5.2 Qualitative Evaluation
We qualitatively asses the effect of the number of Euler steps on generation quality using the CelebA-HQ dataset ($256\times 256$), utilizing the Fréchet Inception Distance (FID) [10.5555/3295222.3295408, Seitzer2020FID] as metric. All FID scores are computed using 50,000 generated samples. As shown in Figure 6, with very few steps (151-5) the samples are blurry but roughly coherent; increasing the steps to 104010-40 sharply improves the quality and semantic detail.
Figure 6: Effect of the number of Euler steps on sample quality. Columns correspond to $1,2,3,4$, $5,10,20,40,80,$ and $250$ Euler steps (left to right). Image sharpness and realism steadily improve as the number of steps increases.
The qualitative analysis shows a strong dependence on the number of Euler steps for CelebA-HQ 256 generation quality. The FID score experiences its most significant reduction initially, dropping sharply from $26.34$ at $20$ steps to $17.12$ at $40$ steps, corresponding to the rapid transition from blurry to sharp samples seen in Figure 7. Beyond $40$ steps, the curve exhibits clear diminishing returns: FID improves from about $12.024$ to only $10.30$ at $250$ steps. Therefore, while $250$ steps yields the best score, near-optimal, computationally efficient generation is already achieved around 8080-$100$ Euler steps.
Figure 7: Effect of the number of Euler steps on sample quality, measured by FID on CelebA-HQ $256$. FID decreases rapidly up to about $40$ steps, reflecting improved sample coherence, and then flattens, with strong diminishing returns beyond roughly $100$ steps.
Table 1: FID (50k samples) on CelebA-HQ-256 ($\times256$).
[…]
Glow [kingma_glow_2018] Flow 68.93
SDE [song2021scorebased] Score-based 7.23
LSGM [vahdat_score-based_2021] Latent score 7.22
LDM-4 [Rombach_2022_CVPR] Latent diffusion 5.11
GAF (ours) Endpoint regression 10.32
Table 1 reports FID scores on CelebA-HQ $256$ for a range of likelihood-based, score-based, and latent-diffusion models. GAF achieves a FID of $10.32$, placing it in a comparable range to established score-based methods on this dataset, while more heavily engineered latent-diffusion models still obtain the best reported numbers. This shows that our endpoint formulation is already competitive, with clear potential for further gains from architectural and training refinements.

### v2

@@ 3.3 Transport Algebra
Transport Algebra organizes GAF’s endpoint pairs into an algebra of endpoint predictors, enabling composable operations that yield controlled data generation. Recall that GAF defines two time-conditioned endpoints $J$ and $K$. Because GAF is modular, the shared trunk can host multiple tuplets $\{(J_{n},K_{n})\}_{n=1}^{N}$. This creates a natural algebra in which compositional operations can be performed (i) directly on the endpoint predictors by combining $J_{n}$ and $K_{n}$ and then applying endpoint refinement, and (ii) by using the emergent velocity field $v_{n}=K_{n}-J_{n}$ and combining these fields algebraically. This ability of GAF to operate algebraically leads to a precise and semantically coherent method of data generation across multiple classes and domains.
@@
GAF’s independent $K$ heads enable compositional transport through endpoint operations (IER) and through linear velocity operations. Let $i,j\in\{1,\dots,N\}$.
@@
    Single-class transport. Samples from class $i$ can be generated by applying IER algorithm using the corresponding $K_{i}$ head to iteratively refine the data endpoint from $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$.
    If velocity view is used and the noise endpoint $J$ is shared, then
@@
    Cross-modality transport (via $J$).
    Move between data endpoints of two (or more) domains. To move from class $i$ to class $j$, we use $J$ as a terminal and $K_{i\to j}$$K_{i}\xrightarrow{J}K_{j}$ (IER) or $v_{i\to j}=v_{i}\xrightarrow{-v_{i}}J\to v_{j}$ (ODE). The $v_{i\to j}$ expands to an encode-decode process:
@@
    This sequential transport process uses $J$ as a universal intermediate representation shared across all classes, enabling transport and composition between any pair of $K$ heads.
@@
    Direct $K$-head interpolation. We can interpolate directly on $K$ heads or in the velocity space via weighted combinations
    $K_{\alpha}=(1-\alpha)K_{i}+\alpha K_{j},$ (28)
    $v_{\alpha}=(1-\alpha)(K_{i}-J)+\alpha(K_{j}-J)=(1-\alpha)v_{i}+\alpha v_{j},$ (29)
@@
    A simultaneous weighted combination of multiple $K$ heads enables compositional generation, which is experimented on in Section 5.6.3. These operations demonstrate GAF’s transport algebra where $K$ independent heads act as linear operators that can be composed, interpolated, and combined to traverse the generative space. This property extends naturally to sequential domains. For example, for video generation, we hypothesize that GAF can smoothly transition between temporal states while preserving semantic coherence through IER interpolation or velocity interpolation, which we will explore in future work.
@@
    Multi-class composition (barycentric). Combine multiple $K$ heads or velocity fields by linear vector operations.
    $K_{\text{blend}}=\sum_{m}w_{m}K_{m}$ (30)
    $v_{\text{blend}}=\sum_{m}w_{m}(K_{m}-J)$ (31)
    where $\sum_{m}w_{m}=1$
By using $J_{n},K_{n}$, IER, $v_{n}$ and the $\{(J_{n},K_{n})\}_{n=1}^{N}$ pairings, together with the swap operator $\mathcal{S}$ (which swaps $J\leftrightarrow K$, $z_{y}\leftrightarrow z_{x}$, and $t\leftrightarrow 1-t$) one can traverse modalities, switch modalities mid generation, revert to the previous modality, chain back to the origin, or even use a specific generated modality as a bootstrap for a chained generation. All of this is possible by switching endpoint heads and applying endpoint refinement, or by switching between the emergent velocity field $v_{i}$, $i\in\{1,\dots,N\}$, and applying simple vector operations. In all cases, $\mathcal{S}$ operator reconfigures bridges and directions on the fly.
@@ 4.1 Network Architecture
An overview of the architecture for image generation is given in Figure 5. We discuss each component below.
Figure 5: GAF architecture for image generation. The DiT-Trunk block follows the architecture of [30]. GAF replaces DiT’s single output projection with two independent heads, $H_{J}$ and $H_{K}$ (see Section 3.1), each producing endpoint residuals that are unpatchified separately and combined with scaled input to form the twins $J$ and $K$.
@@
As shown in Figure 5, we employ a Diffusion Transformer (DiT) [30] architecture as our feature extractor. The GAF trunk takes in either pixel-space representation (e.g., $32\times 32\times 3$ px for low-resolution images) or a latent-space representation (e.g., $32\times 32\times 4$, encoded from high-resolution images such as $256\times 256\times 3$ px using a pretrained variational autoencoder (VAE) [16]). That is because, at higher resolution, pixel-space processing becomes impractical due to memory and compute constraints.
@@
Patchify is the first layer of the trunk, converting the pixel- or latent-space inputs into a sequence of $T$ tokens. It does this by dividing the input in patches, and linearly embedding each patch into the trunk’s hidden dimension $D$. Standard sinusoidal positional embeddings are then added to the tokens to retain spatial information. As noted in [30], the choice of patch size $p$ presents a crucial tradeoff between computational cost and model granularity, as halving $p$ quadruples the GFLOPS with negligible effect on model parameter count. Therefore, we set $p$ as $2$ in our configurations.
DiT Block.
@@
The $J$ head is a single-layer $\text{MLP}_{J}$, which maps trunk features to noise boundary residuals. The MLP expands from $D$ to $2D$ hidden units via a linear layer, and then projects to $p^{2}\times C$ features, where $C$ is the number of channels ($4$ for latents, 3 for RGB pixels). The final layer of $\text{MLP}_{J}$ is initialized to zero for stability. Finally, the outputs are reshaped (i.e., unpatchified) back to their spatial form ($C\times H\times W$).
@@
For each class $k\in\{0,\dots,N-1\}$, each corresponding $K$ head is an independent single-layer $\text{MLP}_{K}$ with identical architecture to $J$. The $K$ heads share no parameters across classes, enabling independent transport operators for each class. During the forward pass, trunk features are computed once for the entire batch, then routed to the appropriate $K$ heads based on class labels. The routing ensures each $K$ head trains on only its assigned class, while the trunk learns shared representations across all classes. Therefore, a single linear projection is sufficient for high-fidelity generation, confirming that the backbone $\Phi$ captures the necessary feature information for the twins.
@@ 4.2 Training
We train GAF with three loss components defined in (24) using mixed-precision (bfloat16) training.
@@
where $\mathcal{L}_{\text{pair}}$ anchors the twins $J$ and $K$ to their endpoints ($\mathbf{z}_{y}$ and $\mathbf{z}_{x}$), $\mathcal{L}_{\text{res}}$ penalizes the residuals near endpoints (as $t\rightarrow 0$ for $J$ and as $t\rightarrow 1$ for $K$), and $\mathcal{L}_{\text{swap}}$ enforces time-reversal symmetry at the midpoint for consistency and self-correction.
@@
4.3 Endpoint Anchoring and Boundary Behavior
[…]
Figure 6: Endpoint anchoring and boundary behavior diagnostics on AFHQ held-out dataset, averaged across classes.
@@ 4.4 Sampling
GAF supports three sampling modes: IER (Section 4.4.1), ODE (Section 4.4.2), and hybrid (Section 4.4.3).
4.4.1 Iterative Endpoint Refinement Sampling
Given the GAF factorization of its endpoint $J$ and $K$, we sample by iteratively refining endpoints on a linear bridge. In the forward direction, $K$ is used to refine the data endpoint from the intermediate bridge points $\mathbf{x_{t}}$. In the reverse direction, $J$ is used to refine the noise endpoint from intermediate bridge points. Both the bridge $\mathbf{x_{t}}=(1-t)\mathbf{z_{0}}+t\mathbf{x_{1}}$ and the endpoints update ($\mathbf{x}_{1}\leftarrow(1-\alpha)\,\mathbf{x}_{1}+\alpha\,K$ for data, $\hat{\mathbf{z}}\leftarrow(1-\alpha)\,\hat{\mathbf{z}}+\alpha\,J$ for noise) are linear operations, so IER traverses straight paths between noise and data without ODE integration.
Forward Generation (noise $\to$ data). Starting from $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$, we initialize $\mathbf{x_{1}}\leftarrow\mathbf{z_{0}}$, set $\alpha\in[0,1]$, and iterate over $t~\in[t_{\varepsilon},1-t_{\varepsilon}]$ (Algorithm 1). At each step, we form $\mathbf{x_{t}}=(1-t)\mathbf{z_{0}}+t\mathbf{x_{1}}$, predict $K$, and update $\mathbf{x_{1}}$ by a convex combination of the current data estimate and the prediction.
Reverse Inversion (data $\to$ noise). Given the data anchor $\mathbf{x_{\text{anchor}}}$, we initialize $\hat{\mathbf{z}}\sim\mathcal{N}(0,I)$, set $\alpha\in[0,1]$, and iterate over $t~\in[t_{\varepsilon},1-t_{\varepsilon}]$ (Algorithm 2). At each step, we form $\mathbf{x_{t}}=(1-t)\hat{\mathbf{z}}+t\mathbf{x_{\text{anchor}}}$, predict $J$, and update $\hat{\mathbf{z}}$ by a convex combination of the current noise estimate and the prediction.
Input: noise $\mathbf{z}_{0}$, steps $S$, $t_{\varepsilon}$, $\alpha$, $\mathrm{GAF}(\cdot,t)$data anchor $\mathbf{x}_{\mathrm{anchor}}$
Output: $\mathbf{x}_{1}$ (data estimate)
$\mathbf{x}_{1}\leftarrow\mathbf{z}_{0}$
$t_{1},\dots,t_{S}\leftarrow\mathrm{Linspace}(t_{\varepsilon},\,1-t_{\varepsilon},\,S)$
for $s\leftarrow 1$ to $S$ do
 $t\leftarrow t_{s}$
 $\mathbf{x}_{t}\leftarrow(1-t)\,\mathbf{z}_{0}+t\,\mathbf{x}_{1}$
 $(\_,K)\leftarrow\mathrm{GAF}(\mathbf{x}_{t},t)$
 $\mathbf{x}_{1}\leftarrow(1-\alpha)\,\mathbf{x}_{1}+\alpha\,K$
end for
return $\mathbf{x}_{1}$
Algorithm 1 IER Forward (noise $\rightarrow$ data; $K$-only)
Input: data anchor $\mathbf{x}_{\mathrm{anchor}}$, steps $S$, $t_{\varepsilon}$, $\alpha$, $\mathrm{GAF}(\cdot,t)$
Output: $\hat{\mathbf{z}}$ (noise estimate)
$\hat{\mathbf{z}}\sim\mathcal{N}(0,I)$
$t_{1},\dots,t_{S}\leftarrow\mathrm{Linspace}(1-t_{\varepsilon},\,t_{\varepsilon},\,S)$
for $s\leftarrow 1$ to $S$ do
 $t\leftarrow t_{s}$
 $\mathbf{x}_{t}\leftarrow(1-t)\,\hat{\mathbf{z}}+t\,\mathbf{x}_{\mathrm{anchor}}$
 $(J,\_)\leftarrow\mathrm{GAF}(\mathbf{x}_{t},t)$
 $\hat{\mathbf{z}}\leftarrow(1-\alpha)\,\hat{\mathbf{z}}+\alpha\,J$
end for
return $\hat{\mathbf{z}}$
Algorithm 2 IER Reverse (data $\rightarrow$ noise; $J$-only)
4.4.2 ODE Sampling
At inference, we generate samples by integrating the emergent velocity field v=KJv=K-J from noise ($t=0$) to data ($t=1$) using Euler’s method [15]. Unlike standard models, GAF defines this field algebraically, enabling precise compositional generation directly within the ODE solver. As during training, we discretize the time interval $t_{\text{eps}},1-t_{\text{eps}}$] into uniformly spaced steps:
$t_{0}=t_{\text{eps}},t_{1},\dots,1-t_{\text{eps}},\text{with}~t_{\text{eps}}=1\times 10^{-3}$
@@ 4.4.2 ODE Sampling
4.4.3 Hybrid Sampling
GAF supports mid-trajectory switching between IER and ODE at any step, in both forward (generation) and reverse (inversion) directions. Run Algorithm 1 until $t_{s}$, then continue with Algorithm 2, or vice versa.
@@ 5 Experimental Evaluation
In this section, we evaluate GAF for image generation, as proposed in Sections 3 and 4. We first describe the dataset (Section 5.1.1) and training setup (Section 5.1.2), then present results for single-class (Section 5.2) and multi-class generation (Section 5.3), followed by qualitative evaluation (Section 5.4), an ablation study (Section 5.5), transport algebra (Section 5.6), and a comparison with Rectified Flow (Section 5.7).
@@ 5.1.1 Dataset
We train and evaluate GAF on four datasets.
CIFAR-10 [18] contains 50,000 training images across 10 object classes at $32\times 32$ px resolution, with $5,000$ images per class. We use this dataset for a $10$-class (i.e., one $J$ and ten $K$) generation task.
CelebA-HQ [25] consists of $202,599$ celebrity face images resized to $256\times 256$ px resolutions. We use this dataset for a single-class (i.e., one $J$ and one $K$) generation task.
AFHQ [5] contains $15,000$ high-quality images of cats, dogs, and wildlife at $512\times 512$ px resolution. We use this dataset for a $3$-class (i.e., one $J$ and three $K$) generation task.
ImageNet-1k [6, 33] is the full ImageNet classification dataset with $1.28$ million training images across $1,000$ object classes, with $1,300$ images per classes. For demonstration purposes, we use a single-class dataset (school bus). Additionally, we use a multi-class subset consisting of $1,000$ distinct classes. For this multi-class generation, we use $1,000$ independent $K$ heads, each corresponding to a specific class.
@@ 5.1.2 Training
Through an extensive parameter sweep of the residual and swap loss coefficients, we found $\lambda_{\text{res}}=0.003$ and $\lambda_{\text{swap}}=0.002$ to perform best. We use these values in all experiments unless stated otherwise. The training images are first encoded by the SD1.5-VAE [32]. We use the AdamW optimizer with $\beta=(0.9,0.99)$, and learning rate $1\times 10^{-4}$ for all of the datasets, with no weight decay in any of the training runs. The ImageNet-1k School Bus class is trained for $300,000$ iterations, whereas the other models are trained for $100,000$ iterations.
For the ImageNet multi-class experiment, we initialize GAF’s trunk from pretrained DiT-XL/2 weights and train the trunk and the $J/K$ heads jointly. We refer to this as retrunking: repurposing an existing pretrained backbone as GAF’s shared trunk. The trunk and endpoint heads are architecturally independent, so any compatible pretrained backbone can be adopted as the starting point for GAF training. The complete training configurations for each dataset is shown in Table 2.
Table 2: Training configurations for each dataset.
CIFAR-10 CelebA-HQ ImageNet
Training scratch scratch retrunked
Architecture DiT-B/2 DiT-XL/2 DiT-XL/2
VAE SD 1.5 SD 1.5
Batch size 256 32 128
Iterations 100k 100k 100k
Learning rate 2e-4 1e-4 1e-4
Warmup steps 500 3000 0
$\beta_{1},\beta_{2}$ 0.9, 0.99 0.9, 0.99 0.9, 0.99
EMA decay 0.9995 0.9999 0.9999
$\lambda_{\text{res}}$ 0.003 0.003 0.0001
$\lambda_{\text{time}}$ 0.002 0.002 0.0001